OpenStudy (anonymous):

Evaluate 23 - 7 ÷ (1/3 + 1/4)

OpenStudy (shane_b):

Original problem:\[\frac{23-7}{\frac{1}{3}+\frac{1}{4}}\]First, simplify both the top and bottom: \[\frac{16}{\frac{7}{12}}\]Now just invert and multiply: \[16*\frac{12}{7}=\frac{192}{7}~or~23\frac{3}{7}\]